Nan and Peg started on trips from New York City and traveled in opposite directions. Nan traveled 10 mph faster than Peg. If they were 450 miles apart after 5 hours, how fast was each person traveling?
Answer by mananth(16949)

Nan and Peg started on trips from New York City and traveled in opposite directions. Nan traveled 10 mph faster than Peg. If they were 450 miles apart after 5 hours, how fast was each person traveling?
...
peg speed = xmph
Nan speed = x+10 mph
...
they travel in opposite directions
so speed is added up
x+x+10=2x+10
distance = 450 miles
..
time = 5 hours
...
450/(2x+10)=5
450=5(2x+10)
450=10x+50
add -50
450-50=10x+50-50
400=10x
x=400/10
x= 40 mph Peg's speed
Nan's speed = 40+10=50 mph
...
m.ananth@hotmail.ca